# A Multi-Task and Multi-Label Classification Model for Implicit Discourse Relation Recognition

### v2

@@ 5.2 Fine-Tuning
LR Decay JS Distance (Multi-Label with MAE) F1-Score (Single-Label with CE)
@@
We fine-tuned our model separately with different learning rates for multi-label IDRR and for single-label IDRR. Table 3 shows the results of fine-tuning our model with the MAE loss in the multi-label setting and with the CE loss in the single-label setting. Under both settings, there is a significant increase in performance when training with a learning rate of $lr=1e^{-5}$ and $lr=5e^{-6}$. For this reason we further experimented training with the linear and cosine annealing decay (Loshchilov and Hutter, 2017) functions for these two learning rates. The linear decay function was set to reduce the learning rate during the first 5 epochs to a final value of half of its original value and the cosine annealing decay function was set to vary between the original learning rate and half of its value for two complete periods during the 10 epochs. Based on the results in Table 3, the warm restart mechanism of the cosine annealing decay function with a learning rate $lr=1e^{-5}$ achieved the best results in the multi-label setting and the linear decay function with a learning rate $lr=5e^{-6}$ achieves the best results in the single-label setting.
@@ 6 Results
Table 4 shows the results of our multi- and single-label models on the test set of DiscoGeM, while Table 5 shows the transfer learning results of our single-label model on the different test splits (see Section 4.2) of the PDTB 3.0. The only difference between the multi- and single-label models is that the first was trained with the MAE loss and the second was trained with the CE loss (see Section 3).
@@ 6.1 Results on DiscoGeM
Models JS Distance (Multi-Label) F1-Score (Single-Label)
@@
Since there is no previous work on full multi-label IDRR, we created a random predictor to use as baseline. The baseline predicts a probability distribution over all the senses at each sense level by emulating the annotation protocol of DiscoGeM. The baseline makes ten single-label predictions following the probability distribution of each sense at each level and the final probability distribution is computed as the average of these ten predictions. In the single-label setting, we compared our results against previous work on single-label IDRR using DiscoGeM (Yung et al., 2022; Costa and Kosseim, 2024).  Yung et al. (2022) trained and tested their model on both DiscoGeM and the PDTB 3.0 on single-label classification only at level-2. When using DiscoGeM they adapt it to the single-label setting by generating ten single-label instances for each discourse relation proportional to its probability distribution in the corpus.  Costa and Kosseim (2024) follow a similar approach to ours in the sense that they also train their model on multi-label discourse relations in DiscoGeM to generate a probability distribution from which they select the majority sense as their predicted single-label.
Table 4 shows that our model is consistently better than the baseline on all metrics and significantly outperforms the current SOTA in single-label level-2 classification (Yung et al., 2022) ($55.99$ vs $23.66$). At level-3, our model achieves results at par with SOTA (Costa and Kosseim, 2024) taking the standard deviation into account ($50.82\pm 1.26$ vs $51.38$). Furthermore, it is important to highlight that the model of Costa and Kosseim (2024) was specifically trained to predict senses at level-3, while our proposed model can generalize across all levels.
@@
 Liu and Strube (2023) -- -- $71.15\pm 0.47$ $54.92\pm 0.81$- --
@@
Table 5: Transfer learning results on different test splits of the PDTB 3.0 in single-label classification (weighted F1-score). The Lin and Ji results were averaged across three different runs with random starts, while the Cross results were averaged across all 12 folds. Values in bold show the best score for each metric.
The results of transfer learning from DiscoGeM to the PDTB 3.0 on the single-label classification are shown in Table 5. Our results are compared against the work of Yung et al. (2022) and the SOTA (Long and Webber, 2022; Liu and Strube, 2023). Both SOTA approaches used a RoBERTa based model and were evaluated using the Ji split (Ji and Eisenstein, 2015) and the cross-validation scheme proposed by Kim et al. (2020).
Table 5 shows that our model has a lower performance than previous work on all metrics for which there was a reference result [^4] We could not find any work on single-label IDDR using the Lin split for the PDTB 3.0 corpus.. However, it is important to note that our model was never exposed to any PDTB 3.0 data in its training. The lower performance in this setting can further be explained by the agreement across corpora shown in Table 1 in Section 4.3, where the majority sense in DiscoGeM matched only on $92(30.5\%)$ of the co-annotated discourse relations in the PDTB 3.0. This shows that a direct transfer learning approach is not possible between both corpora. This also explains the contrasting performances shown in Tables 4 and 5 between our model and the one from Yung et al. (2022), which was trained on the two corpora. Our model performs significantly better when evaluated on DiscoGeM ($55.99$ vs $23.66$), but worst when evaluated on the PDTB 3.0 ($33.11$ vs $38.07$). In the future, after training on DiscoGeM, we could subsequently fine-tune our model on the PDTB 3.0 to improve its performance in this setting.
@@
Table 6: Individual per-class results for each sense in level-1 and level-2 on the test set of DiscoGeM and on different test splits of the PDTB 3.0 in single-label classification (weighted F1-score). Senses marked with -- were not present in the test set. The DiscoGeM Test and the PDTB 3.0 Lin and Ji results were averaged across three different runs with random starts, while the PDTB 3.0 Cross results were averaged across all 12 folds.
7 Analysis
To further evaluate the performance and robustness of our model, we present two additional analyses. First, we conduct a per-class evaluation to assess the ability of the model in classifying each individual sense at level-1 and level-2. Second, we investigate the coherence of the predictions by evaluating how well the predicted labels at level-1 and level-2 align with the hierarchical structure of senses of the PDTB 3.0 framework. For easier interpretation, we consider only single-label results.
7.1 Per-Sense Results
The per-sense performance of the model on the test set of DiscoGeM at level-1 and level-2 in Table 6 follows the distribution of senses in the corpus (see Table 8 in Appendix A). Note that the (Condition, Purpose, Equivalence and Manner) senses never occur as the majority label of a relation in DiscoGeM (see Figure 2 in Section 4.1) and are therefore not present in the test set (see Table 9 in Appendix A). Table 6 shows that our model has difficulties predicting under-represented senses in the corpus, such as Synchronous, Contrast, Similarity and Substitution. One possible solution would be to use data augmentation techniques for these under-represented instances. The confusion matrices in Figures 3 and 4 in Appendix C provide further details into the per-sense performance on the test set of DiscoGeM. The performance on the test sets of the PDTB 3.0 shown in Table 6 is lower with respect to the performance on DiscoGeM, similar to what was shown in Table 4 in Section 6.1 and in Table 5 in Section 6.2. This drop in performance is more pronounced for some senses, such as Temporal at level-1 and Asynchronous and Instantiation at level-2. Conversely, the model performs better at the level-2 Contrast sense. Since the model was not trained on any PDTB 3.0 data, these discrepancies can be explained by inconsistent annotations across both corpora as highlighted in Table 1 in Section 4.3.
@@ 7.2 Sense Coherence across Levels
Table 7: Percentage of the instances that were classified with coherent senses at level-1 and level-2 on the test set of DiscoGeM and on different test splits of the PDTB 3.0 in single-label classification. Senses marked with -- were not present in the test set and senses marked with $n/a$ were never predicted by the model. The DiscoGeM Test and the PDTB 3.0 Lin and Ji results were averaged across three different runs with random starts, while the PDTB 3.0 Cross results were averaged across all 12 folds.
To assess the coherence of the predictions of the model across sense levels, Table 7 shows the percentage of times a sense at level-1 was predicted with a coherent sense at level-2 (and vice-versa) in the test set of DiscoGeM and in the different test sets of the PDTB 3.0. This enables us to examine the extent to which the predictions of the model are consistent with the hierarchical sense structure defined by the PDTB 3.0 framework.
As Table 7 shows, the Contingency sense at level-1 is very often predicted with a coherent level-2 sense (i.e., Cause, Condition or Purpose). Similarly, the level-2 senses Instantiation, Level-of-Detail and Conjunction are often predicted with a coherent level-1 sense. However, other senses, such as Comparison and Expansion at level-1 and Cause at level-2, are more often predicted with a contradicting sense. A possible explanation might be that these senses are inherently hard to distinguish and the model is biased towards the most represented sense. For instance, the level-2 senses Cause and Asynchronous are one of the most often co-annotated pair of senses in the DiscoGeM and Cause appears approximately three times more often than Asynchronous (see Table 8 in Appendix A), the model might be more biased towards Cause when distinguishing between these two senses. One possible method to generate more coherent predictions across sense levels, would be to input the predicted sense distributions of classification head 1 into the other two classification heads (see Figure 1), as well as input the predicted sense distributions of classification head 2 into classification head 3.
@@ 8 Conclusion
In this work we presented a novel approach to one of the most challenging tasks in discourse analysis: IDRR. We trained a multi-task classification model on DiscoGeM to learn both multi-label and single-label representations of discourse relations. Our model establishes a first benchmark on multi-label IDRR and achieves SOTA results in single-label IDRR using DiscoGeM (see Section 6.1). Our single-label classification model was also evaluated on different test splits of the PDTB 3.0 using transfer learning where it achieved lower performance than existing work (see Section 6.2). These results show that direct transfer learning between both corpora still remains a challenge.
Future research should explore the application of data augmentation techniques, such as paraphrasing, to determine whether augmenting under-represented senses in the DiscoGeM could enhance the performance of the model on individual senses (see Section 7.1). Additionally, future research should investigate whether cascading information from higher-level to lower-level classification heads within the model could improve the coherence of its predictions across sense levels (see Section 7.2). Lastly, it would be worthwhile to examine whether a model trained to learn multi-label representations of discourse relations in DiscoGeM could be be further fine-tuned on the PDTB 3.0 to achieve superior performance in single-label classification using its corpus.
@@ 9 Limitations
There are a few limitations to our work. The first limitation comes from the fact that the Cause+Belief sense was not annotated in DiscoGeM. This forced us to adapt the standard set of 14 level-2 senses proposed by Kim et al. (2020) which is used in other works on IDRR using the PDTB 3.0. To overcome this issue and maintain a 14-label set, we proposed to replace the Cause+Belief level-2 sense with the Similarity level-2 sense. However, both senses belong to different level-1 senses. Ideally, we would have replaced the Cause+Belief sense with another sense under the Contingency level-1 sense, but they were all already considered. So, we replaced it with the Similarity sense which was the most frequent sense of the remaining level-2 senses.
The second limitation of our work comes from the choice of pre-trained language model to generate the embeddings of the pair of discourse arguments. Due to computational and time limitations, we could not explore the use of larger language models, such as one of the recently released Gemini 1.5 models (Reid et al., 2024). This could have substantially increased the performance of our models. Nevertheless, we still acknowledge that the extensive experimental analysis conducted in this work would not have been possible without access to a high-performance computing facility. Finally, the use of these powerful resources also entails not negligible carbon footprint which we did not monitor during any of our experiments in this work.
@@ Appendix B Loss Functions
Model Loss JS Distance (Multi-Label) F1-Score (Single-Label)
@@ Appendix C Confusion Matrices
To provide further details into the per-sense performance on the test set of DiscoGeM in Section 7.1, we generated a confusion matrix at level-1 and level-2 for the results of a single run on the test set of DiscoGeM as shown in Figures 3 and 4, respectively. The confusion matrix in Figure 3 shows that, with the exception with Comparison, the most predicted label per true label always aligns with the correct label. Nevertheless, the confusion matrix also shows that the model is more biased towards the most represented senses level-1 senses Expansion and Contingency (see Table 8 in Appendix A). At the level-2, the confusion matrix in Table 4 shows that the model is not able to predict less represented senses in DiscoGeM (see Table 8 in Appendix A), such as Synchronous, Similarity and Substitution. However, for the other senses, the most predicted label per true label often aligns with the correct label.

### v3

@@ 5.2 Fine-Tuning
LR Decay JS Distance
$\searrow$
(Multi-Label with MAE)
F1-Score
$\nearrow$
(Single-Label with CE)
@@
We fine-tuned our model separately with different learning rates for multi-label and single-label IDRR. Table 3 shows the results of fine-tuning our model with the MAE loss in the multi-label setting and with the CE loss in the single-label setting. Under both settings, we observed performance improvements with learning rates of $1\times 10^{-5}$ and $5\times 10^{-6}$. To further enhance model performance with these two learning rates, we also experimented incorporating two decay functions: linear decay and cosine annealing with warm restarts (Loshchilov and Hutter, 2017). For linear decay, the learning rate was gradually reduced over the first 5 epochs to half its initial value. In the cosine annealing schedule, the learning rate oscillated between its original value and half that value over two complete cycles across 10 epochs. As shown in Table 3, the cosine annealing strategy with a learning rate of $1\times 10^{-5}$ achieved the best results for multi-label classification, while the linear decay schedule with a learning rate of $5\times 10^{-6}$ proved most effective in the single-label setting.
@@ 6 Results
In this section, we present the results of our experiments in two parts. Section 6.1 reports the performance of our models on the DiscoGeM test set for both multi-label and single-label IDRR and Section 6.2 presents the transfer learning results of our single-label model evaluated on various test splits of the PDTB 3.0 corpus.
@@ 6.1 Results on DiscoGeM
Models JS Distance
$\searrow$
(Multi-Label)
F1-Score
$\nearrow$
(Single-Label)
@@
Table 4 presents the performance of our models (trained with MAE and CE loss functions) on the DiscoGeM test set for both multi-label and single-label IDRR. Results in the multi-label setting are reported using Jensen-Shannon (JS) distance, while results in the single-label setting are reported using weighted F1-score. As there is currently no prior work on full multi-label IDRR, we include a baseline for comparison. This baseline emulates the DiscoGeM annotation protocol by generating ten single-label predictions per relation, sampled from the probability distribution of each sense at each level, and then averaging them into a probability distribution. In the single-label setting, we compare our results against prior work evaluated on DiscoGeM (Yung et al., 2022; Costa and Kosseim, 2024). As shown in Table 4, our model consistently outperforms the random baseline across all levels and achieves a substantial improvement over the current SOTA in level-2 classification (Yung et al., 2022). At level-3, our model performs at par with the best reported results from Costa and Kosseim (2024). It is worth noting that their model was specifically trained to predict level-3 senses, whereas our model is designed to generalize across all three sense levels simultaneously.
@@
The results of transfer learning from DiscoGeM to PDTB 3.0 on single-label IDRR are presented in Table 5. We compare our model against SOTA approaches in this task (Long and Webber, 2022; Liu and Strube, 2023; Zhao et al., 2023; Zeng et al., 2024; Long and Webber, 2024) and Yung et al. (2022). As shown in Table 5, our model has a lower performance than those trained on PDTB 3.0. However, it is important to emphasize that our model was trained exclusively on DiscoGeM and did not see any PDTB 3.0 data in its training. The lower performance in this zero-shot transfer learning setting, where no fine-tuning is done on the target corpus, can also be explained by thelimited agreement between the two corpora - as shown in Table 1, the majority sense in DiscoGeM only matched the PDTB 3.0 sense on $92$ $(30.5\%)$ of the co-annotated discourse relations. This also explains the contrast in performances shown in Tables 4 and 5 between our model and the one from Yung et al. (2022), which was trained on the two corpora. Our model performs significantly better when evaluated on DiscoGeM, but worst when evaluated on the PDTB 3.0. These findings suggest that, to improve cross-corpus generalization, future work should explore intermediate fine-tuning on PDTB 3.0 after pretraining on DiscoGeM.
@@
Liu and Strube (2023) -- -- $71.15\pm 0.47$ $54.92\pm 0.81$ $70.06\pm 1.72$ $55.26\pm 1.32$
Zhao et al. (2023) -- -- $69.06\pm n/a$ $52.73\pm n/a$- --
Zeng et al. (2024) -- -- $71.59\pm n/a$ $56.50\pm n/a$- --
Long and Webber (2024) -- -- $71.19\pm n/a$ $52.91\pm n/a$- --
@@
Table 5: Transfer learning results on different test splits of the PDTB 3.0 in single-label classification (weighted F1-score). The Lin and Ji results were averaged across three different runs with random starts, while the Cross results were averaged across all 12 folds. Values in bold show the best score for each level.
7 Analysis
To further evaluate the performance of our model, we conduct two additional analyses. In Section 7.1, we perform a per-sense evaluation to assess the ability of the model to correctly classify individual senses at both level-1 and level-2. In Section 7.2, we evaluate the cross-level coherence of the predictions by examining how well the predicted labels at level-1 and level-2 align with the sense hierarchy defined in the PDTB 3.0 framework. For ease of interpretation, we consider only single-label results.
7.1 Per-Sense Results
@@
Table 6: Individual per-sense results for each sense in level-1 and level-2 on the test set of DiscoGeM and on different test splits of the PDTB 3.0 in single-label classification (weighted F1-score). Senses marked with "--" were not present in the DiscoGeM test set. The DiscoGeM and the PDTB 3.0 Lin and Ji results were averaged across three different runs with random starts, while the PDTB 3.0 Cross results were averaged across all 12 folds.
As shown in Table 6, the per-sense performance of the model on the DiscoGeM test set at both level-1 and level-2 generally reflects the distribution of senses in the corpus (see Table 8 in Appendix A). Note that some certain senses — Condition, Purpose, Equivalence and Manner — never appear as the majority label of relation in DiscoGeM and are therefore absent from the test set (see Figure 2 and Table 9 in Appendix A). Unsurprisingly, the model struggles to predict under-represented senses such as Synchronous, Contrast, Similarity and Substitution, which appear infrequently in the training data. Addressing this imbalance, potentially through targeted data augmentation, could improve generalization on these under-represented senses. The confusion matrices in Figures 3 and 4 in Appendix C provide further details into the per-sense performance on the test set of DiscoGeM. The performance of the model on the PDTB 3.0 test sets, also shown in Table 6, is consistently lower than on DiscoGeM, mirroring the trends reported in Table 4 and Table 5. This performance gap is particularly pronounced for certain senses, such as Temporal at level-1 and Asynchronous and Instantiation at level-2. Conversely, the model performs better at the level-2 Contrast sense. Since the model was not trained on any PDTB 3.0 data, these discrepancies can be explained by annotation inconsistencies across both corpora as highlighted in Section 4.3.
@@ 7.2 Sense Coherence across Levels
Table 7: Percentage of the instances that were classified with coherent senses at level-1 and level-2 on the test set of DiscoGeM and on different test splits of the PDTB 3.0 in single-label classification. Senses marked with "--" were not present in the DiscoGeM test set and senses marked with "$n/a$" were never predicted by the model. The DiscoGeM and the PDTB 3.0 Lin and Ji results were averaged across three different runs with random starts, while the PDTB 3.0 Cross results were averaged across all 12 folds.
Table 7 shows the percentage of times a sense at level-1 was predicted with a coherent sense at level-2 (and vice-versa) in the test set of DiscoGeM and in the different test sets of the PDTB 3.0. This enables us to examine the extent to which the predictions of the model are consistent with the hierarchical sense structure defined in the PDTB 3.0.
As Table 7 shows, the Contingency sense at level-1 is very often predicted with a coherent level-2 sense (i.e., Cause, Condition or Purpose). Similarly, the level-2 senses Instantiation, Level-of-Detail and Conjunction are often predicted with a coherent level-1 sense. However, other senses, such as Comparison and Expansion at level-1 and Cause at level-2, are more often predicted with a contradicting sense. A possible explanation might be that these senses are inherently hard to distinguish and the model is biased towards the most represented sense. For instance, the level-2 senses Cause and Asynchronous are among the most often co-annotated pair of senses in the DiscoGeM and Cause appears approximately three times more often than Asynchronous (see Table 8 in Appendix A), the model is likely more biased towards Cause when distinguishing between these two senses. One possible method to generate more coherent predictions across sense levels, would be to share the predictions of lower-level senses to help inform the prediction of higher-level senses within the model.
@@ 8 Conclusion
In this work, we proposed a novel multi-label framework for IDRR, addressing one of the most complex challenges in discourse analysis. We trained a multi-task classification model on the DiscoGeM corpus to simultaneously learn multi-label representations of discourse relations across all three sense levels in the PDTB 3.0 framework. Our model can also be adapted to the traditional single-label IDRR setting by selecting the sense with the highest probability in the multi-label representation. We conducted extensive experiments to identify optimal model configurations and loss functions in both settings. Our approach establishes a first benchmark on multi-label IDRR and achieves SOTA results in single-label IDRR using DiscoGeM. Finally, we evaluated our model on the PDTB 3.0 corpus in the single-label setting, presenting the first analysis of transfer learning between the DiscoGeM and PDTB 3.0 corpora for IDRR. Our results show that zero-shot direct transfer learning between both corpora still needs further research.
Future work should explore the application of data augmentation techniques, such as paraphrasing, to determine whether augmenting under-represented senses in the DiscoGeM could enhance the performance of the model on individual senses. Additionally, future research should investigate whether cascading information from higher-level to lower-level classification heads within the model could improve the coherence of its predictions across sense levels. Lastly, it would be worthwhile to examine whether a model trained to learn multi-label representations of discourse relations in DiscoGeM could be further fine-tuned on the PDTB 3.0 to achieve superior performance in single-label IDRR using the PDTB 3.0 corpus.
@@ 9 Limitations
Despite the promising results, there are a few limitations to our work. The first limitation of our work comes from the fact that the CAUSE+BELIEF sense was not annotated in the DiscoGeM corpus. Therefore, we could not use the standard 14-label set of second-level senses proposed for the PDTB 3.0 by Kim et al. (2020) and widely used in literature. Instead, we replaced it by the SIMILARITY sense - the next most available sense in the DiscoGeM corpus. Ideally, we would have replaced the Cause+Belief sense with another sense under the Contingency level-1 sense, but they were all already considered.
Our choice of pre-trained language model to generate the embeddings for each pair of discourse arguments also imposes a limitation to our work. Due to computational and time limitations, we could not explore the fine-tuning of larger models, such as LLaMA 3 (Grattafiori et al., 2024), nor explore prompting LLMs. However, in our other work (Costa and Kosseim, 2025), we show that the use of LLMs via direct prompting with few-shot learning for IDRR does not lead to better results. Finally, we would like to acknowledge that the extensive experimental analysis conducted in this work would not have been possible without access to a high-performance computing facility - which entails a non-negligible carbon footprint which we did not monitor during any of our experiments.
Acknowledgements
The authors would like to thank the anonymous reviewers for their comments. This work was financially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC).
@@ Appendix A Data Statistics
Figure 2: Distribution of majority label senses across all splits of the DiscoGeM corpus (see Section 4.1).
@@ Appendix B Loss Functions
Model Loss JS Distance
$\searrow$
(Multi-Label)
F1-Score
$\nearrow$
(Single-Label)
@@ Appendix C Confusion Matrices
To provide further details into the per-sense performance on the test set of DiscoGeM in Section 7.1, we generated a confusion matrix at level-1 and level-2 for the results of a single run on the test set of DiscoGeM as shown in Figures 3 and 4, respectively. The confusion matrix in Figure 3 shows that, with the exception with Comparison, the most predicted label always aligns with the correct label. Nevertheless, the confusion matrix also shows that the model is more biased towards the most represented senses level-1 senses Expansion and Contingency (see Table 8 in Appendix A). At the level-2, the confusion matrix in Table 4 shows that the model is not able to predict less represented senses in DiscoGeM (see Table 8 in Appendix A), such as Synchronous, Similarity and Substitution. However, for the other senses, the most predicted label often aligns with the correct label.